# LLM-HDR: Bridging LLM-based Perception and Self-Supervision for Unpaired LDR-to-HDR Image Reconstruction

### v2

@@ 1 Introduction
To address this gap, we propose LLM-HDR, a method that leverages LLM-driven cycle consistency [117, 17] objective using unpaired data, where the model learns a mapping between domains, i.e., LDR $\leftrightarrow$ HDR. In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing Contrastive Language-Image Pretraining (CLIP) [81] embeddings and loss based on semantic segmentation. Moreover, LLM-based artifact- and exposure-aware information further supports the training process to deliver more realistic and natural HDR image generation. The proposed method reconstructs artifact-free and visually impressive HDR from single-exposed LDR images. It also outperforms the most recent state-of-the-art which predominantly uses paired datasets for training. Our work makes the following contributions:
@@
  • We leverage LLM-based loss and artifact- and exposure-aware saliency maps using human-defined prompting to further refine those areas in the reconstructed HDR images (Secs. 3.1 and 3.2).
@@ 3.1 Architecture Modules
LLM. We adopt a pre-trained LLM [38] for zero-shot Q&A sessions (using human-defined knowledge) in each iteration of the training process. We prompt the LLM to provide the pixels identified as artifacts, and over/underexposed regions for both the reconstructed HDR and LDR. Given that these three factors constitute the nature of the LDR $\leftrightarrow$ HDR translation, the pixels proposed by the LLM are used in a LLM-based loss function. In addition, the LLM generates a separate saliency map for each of those areas, i.e., artifact, overexposed, and underexposed maps. The artifact map is multiplied with the input features to the feedback mechanism of the generators $G_{Y}$ and $G_{X}$. The underexposed map is fed into each of the skip connections (levels $1$ to $3$) in the generator $G_{Y}$ only and the overexposed map is multiplied with the input features to the bottleneck layer of this generator. We use element-wise multiplication as a gating mechanism to emphasize the selected areas in the maps. The rationale behind the separated saliency map fusion is based on the proposed U-Net architecture where distorted pixels are mostly refined by the feedback mechanism and later layers, high-intensity lighting details are generally recovered by the middle or later layers, and low-intensity shadows are reconstructed realistically at the starting and ending layers (Sec. S2.2 for visual depiction of the proposed LLM module and implementation details).
@@ 3.2 Loss Functions
LLM Loss. Given a reconstructed HDR/LDR image, this loss is based on pixels identified by the LLM as artifacts and over/underexposed regions during Q&A sessions (using human-defined knowledge). The loss is defined as:
$\mathcal{L}_{\text{llm}}=3\times\frac{\hat{y_{\text{af}}}}{\hat{y_{\text{total}}}}+2\times\frac{\hat{y_{\text{ox}}}}{\hat{y_{\text{total}}}}+1.5\times\frac{\hat{y_{\text{ux}}}}{\hat{y_{\text{total}}}},$ (2)
where $\hat{y_{\text{total}}}$ represents the total number of pixels in the reconstructed HDR $\hat{y}$, while $\hat{y_{\text{af}}}$, $\hat{y_{\text{ox}}}$, and $\hat{y_{\text{ux}}}$ denote the total number of pixels in the areas perceived by the LLM as artifacts, overexposed, and underexposed. The weights on each of the three components are selected after thorough ablations as well as theoretical considerations. Our analysis suggests that artifact mitigation is more important than exposure issues for the perceptual quality of the reconstructed images because artifacts make the images unnatural and unpleasing to the human eye. Humans have greater tolerance towards dark areas than faded bright regions hence we give lower weight to the third component. (Sec. S2.3 for visual depiction of the LLM loss). This loss is calculated in the backward cycle with $\hat{x}$ but only with the artifact component. The rationale behind this decision lies in the nature of the output, i.e., LDR can be overexposed or underexposed.
Contrastive Loss. This loss is based on embeddings extracted using a CLIP encoder and ensures semantic information preservation across domains. Here, we do not directly extract the embeddings from LDR images but instead, use a histogram-equalized version processed using the OpenCV function equalizeHist [10]. Histogram equalization improves the pixel visibility in extreme lighting areas or areas with shadows/darkness of an image by re-adjusting the contrast and saturation levels. This is done by spreading out the frequent pixel intensity values across $256$ bins. Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image. For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(\hat{y})$ we first define the cosine similarity between them as follows:
[…]
$\begin{split}\mathcal{L}_{\text{con}}=-\frac{1}{N}\times\\ \sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\bar{x}_{i},\bar{y}_{i})/\tau)}{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(\exp(\text{sim}(\bar{x}_{i},\bar{x}_{j})/\tau)+\exp(\text{sim}(\bar{x}_{i},\bar{y}_{j})/\tau))},\end{split}$ (4)
@@
Semantic Segmentation Loss. This loss is based on segmentation masks. The Mean Intersection over Union (mIoU) metric measures the amount of overlap between ground truth and predicted segmentation masks. Similar to the previous loss function, we use histogram-equalized versions of the LDR images processed using the OpenCV function equalizeHist [10]. We choose equalized images instead of original LDR because segmentation of low-light or extremely bright images does not yield good results. We use Segment Anything (SAM) [50] to generate segmentation classes for the histogram-equalized LDR and reconstructed tone-mapped HDR images (Sec. S2.3 for visual depiction of the semantic segmentation loss). This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images. We can define the IoU metric as:
$\text{IoU}_{c}=\frac{x_{c}\cap\hat{y_{c}}}{x_{c}\cup\hat{y_{c}}},$ (5)
where $x_{c}$ and $\hat{y_{c}}$ represent the segmentation for class $c$ in image $x$ and $\hat{y}$, respectively. $\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$. The mean IoU over all segmentation classes can be formulated as:
$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (6)
@@ 5.1 HDR Reconstruction
Paired Datasets. Tab. 2 summarizes the results of evaluation on HDRTV [16], NTIRE [79], HDR-Synth & HDR-Real [64] paired datasets, the most widely used datasets by the state-of-the-art (Sec. S3.1 for an extensive qualitative evaluation). In the top part of the table, we compare our method with $11$ supervised learning approaches. The proposed method outperforms those approaches in terms of SSIM, LPIPS, and HDR-VDP-2. In terms of PSNR, it is second best after Ghost-free HDR. In the bottom part of the table, we compare our method with $5$ unsupervised, weakly-supervised, and self-supervised learning methods. The proposed method outperforms those approaches on all evaluation metrics. It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training. However, in all experiments, our method is trained in a self-supervised fashion with unpaired samples (Sec. 3).
@@ 6 Conclusion
Figure S3: Overview of the proposed LLM module. Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not. Must response with 1) Yes or No only, 2) If Yes, can you get me the saliency maps of the artifacts (in red), overexposed (in blue) and underexposed (in yellow) areas of this image?”, System: “Outputs the saliency maps for the detected artifacts in red, overexposed pixels in blue, and underexposed pixels in yellow.”
@@
Figure S4: Depiction of the LLM-based loss $\mathcal{L}_{\text{llm}}$. Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not. Must response with 1) Yes or No only, 2) If Yes, return the [artifact areas number of pixels, total number of pixels]. Also highlight the artifact areas with pink. Similarly find if there are any over/under exposure areas in the scene and return the [pixels in overexposure, pixels in underexposure areas] and highlight the overexposed areas in blue and underexposed areas in green.”, System: “Outputs the number of pixels with the pixel list (in .txt files) for the detected artifacts in pink, overexposed pixels in blue, and underexposed pixels in green.”
@@
LLM. The artifact-aware saliency map $LLM_{\text{af}}$ is extracted as feature of size $512$ with $1\times 1$ Conv and we perform an element-wise multiplication with the output features of the bottleneck layer (i.e., layer $4$) of the U-Net generators $G_{Y}$ and $G_{X}$ which goes into the feedback mechanism. The exposure-aware saliency map $LLM_{\text{ox}}$ (for overexposure) (which goes into the generator $G_{Y}$ only) is extracted as feature of size $256$ with $1\times 1$ Conv and fused with the output features of layer $3$ henceforth concatenated with the encoder embeddings which goes into the bottleneck layer. Finally, $LLM_{\text{ux}}$ (for underexposure) is extracted as feature of sizes $64$, $128$, and $256$, with $1\times 1$ Conv and fused with the skip connections in each level of the U-Net generator $G_{Y}$ (i.e., level 1, 2, and 3). Fig. S3 depicts the overview of the proposed $LLM$ module.
@@
Figs. S4, S5, S6 and S7 illustrate the formulation of the LLM-based loss, contrastive, semantic segmentation, and cycle consistency functions, respectively.
@@
Figs. S12, S13, S14, S15, S16, S17 and S18 illustrate the quality of the HDR images reconstructed with our method. Our method mitigates artifacts such that the resulting HDR closely resembles the ground truth HDR. The finer textures, color hues, and shades are well preserved compared to the other methods. The granular level details in the bright and dark regions are also well reconstructed.
@@
Table S3: Architecture ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets. Uo: The U-Net generator used in SingleHDR(W) [55], Uf: The proposed U-Net generator, E: CLIP embedding encoder, LLM: LLM module for artifact-aware and exposure-aware maps.
@@
Architecture. Our method uses the cycle consistency concept of CycleGAN and introduces many new components to address the LDR $\leftrightarrow$ HDR translation task. The most significant components are the LLM module, artifact-aware feedback and exposure-aware skip connection-based U-Net generators, and the CLIP embedding encoders (Sec. 3.1). We perform a thorough ablation on each of the components to highlight the contributions. Tab. S3 summarizes the results where we add each component to the architecture as we go down. The first row lists the results from the original CycleGAN method. The second row is the CycleGAN utilizing the U-Net generator from SingleHDR(W). The third row is the CycleGAN utilizing our feedback U-Net generator. This results in a significant improvement in SSIM, LPIPS, and HDR-VDP-2 (Q-Score). The fourth row shows the results when we add the LLM module to the architecture which shows some improvement in the PSNR, LPIPS, and Q-Score. The fifth row lists the results when we remove the LLM module and add the CLIP embedding encoders to the architecture. This results in an improvement in PSNR, SSIM, and Q-Score. The sixth row depicts good improvement in almost all the metrics as we add both CLIP encoders and LLM module to the architecture. The seventh, eight, and ninth rows depict the improvements after adding contrastive, semantic segmentation, and LLM-based objectives (Sec. 3.2) in the architecture, respectively. The tenth row shows the improvements when we add only the contrastive and semantic segmentation objectives in the architecture. The eleventh row depicts the results with all components in our architecture but with the original U-Net from SingleHDR(W). The last row is the full architecture with all components which achieves the best performance on all metrics.
Loss Functions. We also study the loss functions and their contribution towards the overall results keeping the architectural components constant. Tab. S4 summarizes the results. The first row is the results from the standard adversarial and cycle consistency loss. In the second row we add the identity loss which improves all metrics. In the third, fourth, and fifth rows we add the contrastive, semantic segmentation, and LLM-based objectives. This results in an improvement in all metrics. The sixth row shows the results when we add contrastive and semantic losses together which gives significant improvement in the metrics. Finally, adding all loss functions (seventh row) achieves the best results on all metrics.
@@
Finally, we experiment with the weights for the loss components with respect to PSNR. We perform tests on the temperature parameter $\tau$ of the contrastive loss and the weight parameters $\alpha$ and $\beta$ for the loss components $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{sem}}$. Fig. S8 shows the results. The temperature parameter $\tau$ gives the optimal PSNR with $\tau=0.08$ while the result degrades if we increase the value further. If we decrease the value, it gives a sub-optimal PSNR at $\tau=0.07$, so we fix the value to $\tau=0.08$ in all experiments. The values of $\alpha$ and $\beta$ give optimal PSNR for $\alpha=2$ and $\beta=2$. The weights for the LLM-based loss are set to $3$, $2$, and $1.5$ for the artifact, overexposed, and underexposed components, respectively. We conducted ablation experiments on these weights considering values from $1$ to $5$ and found that the selected combination of weights yield the best PSNR.

### v3

@@ 1 Introduction
To address this gap, we propose LLM-HDR, a method that leverages LLM-driven cycle consistency [117, 17] objective using unpaired data, where the model learns a mapping between domains, i.e., LDR $\leftrightarrow$ HDR. In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing Contrastive Language-Image Pretraining (CLIP) [81] embeddings and loss based on semantic segmentation. Moreover, LLM-based reasoning of artifact- and exposure-aware information further supports the training process to deliver more realistic and natural HDR image generation. The proposed method reconstructs artifact-free and visually impressive HDR from single-exposed LDR images. It also outperforms the most recent state-of-the-art which predominantly uses paired datasets for training. Our work makes the following contributions:
@@
  • We leverage LLM-based reasoning to obtain a loss and artifact- and exposure-aware saliency maps using human-defined prompting to further refine those areas in the reconstructed HDR images (Secs. 3.1 and 3.2).
@@ 3.1 Architecture Modules
LLM. We adopt a pre-trained LLM [38] for zero-shot Q&A sessions (using human-defined knowledge) in each epoch of the training process as early training outputs are noisy. Evaluating once in every epoch allows the LLM to assess relatively more stable and informative outputs reducing noise in its feedback. We prompt the LLM to provide the reasoning for whether the generated image has artifacts, and over/underexposed regions for both the reconstructed HDR and LDR. Given that these three factors constitute the nature of the LDR $\leftrightarrow$ HDR translation, the reasoning of the images proposed by the LLM along with a heuristic based post-processing using OpenCV [10] to obtain the corresponding pixels are used in a LLM-based loss function. In addition, the LLM’s reasoning also gives way to the generation of a separate saliency map for each of those areas, i.e., artifact, overexposed, and underexposed maps using the same heuristic based post-processing. The artifact map is multiplied with the input features to the feedback mechanism of the generators $G_{Y}$ and $G_{X}$. The underexposed map is fed into each of the skip connections (levels $1$ to $3$) in the generator $G_{Y}$ only and the overexposed map is multiplied with the input features to the bottleneck layer of this generator. We use element-wise multiplication as a gating mechanism to emphasize the selected areas in the maps. The rationale behind the separated saliency map fusion is based on the proposed U-Net architecture where distorted pixels are mostly refined by the feedback mechanism and later layers, high-intensity lighting details are generally recovered by the middle or later layers, and low-intensity shadows are reconstructed realistically at the starting and ending layers (Sec. S2.2 for visual depiction of the proposed LLM module and implementation details).
@@ 3.2 Loss Functions
LLM Reasoning Inspired Loss. Given a reconstructed HDR/LDR image, this loss is based on pixels identified as artifacts and over/underexposed regions by the heuristic based post-processing using OpenCV [10] inspired by the LLM’s reasoning during Q&A sessions (using human-defined knowledge). The loss is defined as:
$\mathcal{L}_{\text{llm}}=\delta_{1}\times\frac{\hat{y_{\text{af}}}}{\hat{y_{\text{total}}}}+\delta_{2}\times\frac{\hat{y_{\text{ox}}}}{\hat{y_{\text{total}}}}+\delta_{3}\times\frac{\hat{y_{\text{ux}}}}{\hat{y_{\text{total}}}},$ (2)
where $\hat{y_{\text{total}}}$ represents the total number of pixels in the reconstructed HDR $\hat{y}$, while $\hat{y_{\text{af}}}$, $\hat{y_{\text{ox}}}$, and $\hat{y_{\text{ux}}}$ denote the total number of pixels in the areas perceived by the LLM as artifacts, overexposed, and underexposed and generated by the heuristic based post-processing. The weights on each of the three components are selected after thorough ablations as well as theoretical considerations. Our analysis suggests that artifact mitigation is more important than exposure issues for the perceptual quality of the reconstructed images because artifacts make the images unnatural and unpleasing to the human eye hence, $\delta_{1}$ is set to 3. Humans have greater tolerance towards dark areas than faded bright regions hence we give lower weight to the third component, i.e., $\delta_{2}$ and $\delta_{3}$ are set to 2 and 1.5 correspondingly. (Sec. S2.3 for visual depiction of the LLM loss). This loss is calculated in the backward cycle with $\hat{x}$ but only with the artifact component. The rationale behind this decision lies in the nature of the output, i.e., LDR can be overexposed or underexposed.
Contrastive Loss. This loss is based on embeddings extracted using a CLIP encoder and ensures semantic information preservation across domains. Here, we do not directly extract the embeddings from LDR images but instead, use a histogram-equalized version processed using the OpenCV function equalizeHist [10]. Histogram equalization improves the pixel visibility in extreme lighting areas or areas with shadows/darkness of an image by re-adjusting the contrast and saturation levels. This is done by spreading out the frequent pixel intensity values across $256$ bins. Equalization often leads to revealing hidden semantic information or non-perceivable objects in an image. For the image embedding $\bar{x}$ from $E(x)$ and $\bar{y}$ from $E(\hat{y})$ we first define the cosine similarity between them as $\text{sim}(\bar{x},\bar{y})$ (Sec. S2.3 for full equation).
We formulate the contrastive loss for an input batch as positive pairs of images (e.g., LDR and the corresponding reconstructed HDR) and negative pairs of images (e.g., each LDR with the rest of the LDR as well as the rest of the reconstructed HDR in the batch) as follows:
$\begin{split}\mathcal{L}_{\text{con}}=-\frac{1}{N}\times\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\bar{x}_{i},\bar{y}_{i})/\tau)}{\sum\limits_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(\exp(\text{sim}(\bar{x}_{i},\bar{x}_{j})/\tau)+\exp(\text{sim}(\bar{x}_{i},\bar{y}_{j})/\tau))},\end{split}$ (3)
@@
Semantic Segmentation Loss. This loss is based on segmentation masks. The Mean Intersection over Union (mIoU) metric measures the amount of overlap between ground truth and predicted segmentation masks. Similar to the previous loss function, we use histogram-equalized versions of the LDR images processed using the OpenCV function equalizeHist [10]. We choose equalized images instead of original LDR because segmentation of low-light or extremely bright images does not yield good results. We use Segment Anything (SAM) [50] to generate segmentation classes for the histogram-equalized LDR and reconstructed tone-mapped HDR images (Sec. S2.3 for visual depiction of the semantic segmentation loss). This loss component helps in mitigating differences in boundary and edge pixels between the LDR and HDR images. We can define the IoU metric as $\text{IoU}_{c}$. The mean IoU over all segmentation classes can be represented as mIoU (Sec. S2.3 for full equations).
@@ 5.1 HDR Reconstruction
Figure 3: Examples of HDR images reconstructed with our method and recent state-of-the-art.
Paired Datasets. Tab. 2 summarizes the results of evaluation on HDRTV [16], NTIRE [79], HDR-Synth & HDR-Real [64] paired datasets, the most widely used datasets by the state-of-the-art (Sec. S3.1 for an extensive qualitative evaluation). In the top part of the table, we compare our method with $11$ supervised learning approaches. The proposed method outperforms those approaches in terms of SSIM, LPIPS, and HDR-VDP-2. In terms of PSNR, it is second best after Ghost-free HDR. In the bottom part of the table, we compare our method with $5$ unsupervised, weakly-supervised, and self-supervised learning methods. The proposed method outperforms those approaches on all evaluation metrics. Although the improvement is small for PSNR and SSIM but LPIPS and Q-Socre show considerable improvement, which is important to note as perceptual, semantic, and high dynamic qualities are captured by these two metrices only. It is worth noting that we used paired datasets in order to compare our method with the state-of-the-art, which predominantly requires paired samples for training. However, in all experiments, our method is trained in a self-supervised fashion with unpaired samples (Sec. 3).
@@ 6 Conclusion
Figure S3: Overview of the proposed LLM module. Here LLM signifies the LLM based reasoning of the input image and the heuristic based post-processing with OpenCV [10] for generation of the saliency maps. Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not. Must response with 1) Yes or No only, 2) If Yes, can you get me the saliency maps of the artifacts (in red), overexposed (in blue) and underexposed (in yellow) areas of this image?”, System: “Outputs the saliency maps for the detected artifacts in red, overexposed pixels in blue, and underexposed pixels in yellow.”
@@
Figure S4: Depiction of the LLM Reasoning Inspired loss $\mathcal{L}_{\text{llm}}$. Here LLM signifies the LLM based reasoning of the input image and the heuristic based post-processing with OpenCV [10] for generation of the affected pixel positions. Q&A sessions include: User prompt: “Tell me if there are any synthesis artifacts in the given scene or not. Must response with 1) Yes or No only, 2) If Yes, return the [artifact areas number of pixels, total number of pixels]. Also highlight the artifact areas with pink. Similarly find if there are any over/under exposure areas in the scene and return the [pixels in overexposure, pixels in underexposure areas] and highlight the overexposed areas in blue and underexposed areas in green.”, System: “Outputs the number of pixels with the pixel list (in .txt files) for the detected artifacts in pink, overexposed pixels in blue, and underexposed pixels in green.”
@@
LLM. We use the reasoning from the LLM to determine whether there is any exposure/artifact issues in the generation. Henceforth, we use heuristic based OpenCV [10] functions to threshold the pixels to generate the exposure-aware masks. Moreover, we also use the Laplacian and GaussianBlur functions of OpenCV to generate the artifact-aware masks. The artifact-aware saliency map $LLM_{\text{af}}$ is extracted as feature of size $512$ with $1\times 1$ Conv and we perform an element-wise multiplication with the output features of the bottleneck layer (i.e., layer $4$) of the U-Net generators $G_{Y}$ and $G_{X}$ which goes into the feedback mechanism. The exposure-aware saliency map $LLM_{\text{ox}}$ (for overexposure) (which goes into the generator $G_{Y}$ only) is extracted as feature of size $256$ with $1\times 1$ Conv and fused with the output features of layer $3$ henceforth concatenated with the encoder embeddings which goes into the bottleneck layer. Finally, $LLM_{\text{ux}}$ (for underexposure) is extracted as feature of sizes $64$, $128$, and $256$, with $1\times 1$ Conv and fused with the skip connections in each level of the U-Net generator $G_{Y}$ (i.e., level 1, 2, and 3). Fig. S3 depicts the overview of the proposed $LLM$ module.
@@
Figs. S4, S5, S6 and S7 illustrate the formulation of the LLM Reasoning Inspired loss, contrastive, semantic segmentation, and cycle consistency functions, respectively.
The $\text{IoU}_{c}$ can be formulated as:
$\text{IoU}_{c}=\frac{x_{c}\cap\hat{y_{c}}}{x_{c}\cup\hat{y_{c}}},$ (10)
where $x_{c}$ and $\hat{y_{c}}$ represent the segmentation for class $c$ in image $x$ and $\hat{y}$, respectively. $\cap$ represents the overlapping area of predicted and ground truth pixels while $\cup$ represents the total area covered by predicted and ground truth pixels for class $c$.
The $\text{mIoU}_{c}$ can be formulated as:
$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\text{IoU}_{c}.$ (11)
The cosine similarity $\text{sim}(\bar{x},\bar{y})$ can be formulated as:
[…]
@@
Figs. S12, S13, S14, S15, S16 and S17 illustrate the quality of the HDR images reconstructed with our method. Our method mitigates artifacts such that the resulting HDR closely resembles the ground truth HDR. The finer textures, color hues, and shades are well preserved compared to the other methods. The granular level details in the bright and dark regions are also well reconstructed.
@@
Table S3: Architecture ablation results with the HDRTV [16], NTIRE [79] and HDR-Synth & HDR-Real [64] datasets. Uo: The U-Net generator used in SingleHDR(W) [55], Uf: The proposed U-Net generator, E: CLIP embedding encoder, LLM: LLM module for artifact-aware and exposure-aware reasoning.
@@
Architecture. Our method uses the cycle consistency concept of CycleGAN and introduces many new components to address the LDR $\leftrightarrow$ HDR translation task. The most significant components are the LLM module, artifact-aware feedback and exposure-aware skip connection-based U-Net generators, and the CLIP embedding encoders (Sec. 3.1). We perform a thorough ablation on each of the components to highlight the contributions. Tab. S3 summarizes the results where we add each component to the architecture as we go down. The first row lists the results from the original CycleGAN method. The second row is the CycleGAN utilizing the U-Net generator from SingleHDR(W). The third row is the CycleGAN utilizing our feedback U-Net generator. This results in a significant improvement in SSIM, LPIPS, and HDR-VDP-2 (Q-Score). The fourth row shows the results when we add the LLM module to the architecture which shows some improvement in the PSNR, LPIPS, and Q-Score. The fifth row lists the results when we remove the LLM module and add the CLIP embedding encoders to the architecture. This results in an improvement in PSNR, SSIM, and Q-Score. The sixth row depicts good improvement in almost all the metrics as we add both CLIP encoders and LLM module to the architecture. The seventh, eight, and ninth rows depict the improvements after adding contrastive, semantic segmentation, and LLM Reasoning Inspired objectives (Sec. 3.2) in the architecture, respectively. The tenth row shows the improvements when we add only the contrastive and semantic segmentation objectives in the architecture. The eleventh row depicts the results with all components in our architecture but with the original U-Net from SingleHDR(W). The last row is the full architecture with all components which achieves the best performance on all metrics.
Loss Functions. We also study the loss functions and their contribution towards the overall results keeping the architectural components constant. Tab. S4 summarizes the results. The first row is the results from the standard adversarial and cycle consistency loss. In the second row we add the identity loss which improves all metrics. In the third, fourth, and fifth rows we add the contrastive, semantic segmentation, and LLM Reasoning Inspired objectives. This results in an improvement in all metrics. The sixth row shows the results when we add contrastive and semantic losses together which gives significant improvement in the metrics. Finally, adding all loss functions (seventh row) achieves the best results on all metrics.
@@
Finally, we experiment with the weights for the loss components with respect to PSNR. We perform tests on the temperature parameter $\tau$ of the contrastive loss and the weight parameters $\alpha$ and $\beta$ for the loss components $\mathcal{L}_{\text{con}}$ and $\mathcal{L}_{\text{sem}}$. Fig. S8 shows the results. The temperature parameter $\tau$ gives the optimal PSNR with $\tau=0.08$ while the result degrades if we increase the value further. If we decrease the value, it gives a sub-optimal PSNR at $\tau=0.07$, so we fix the value to $\tau=0.08$ in all experiments. The values of $\alpha$ and $\beta$ give optimal PSNR for $\alpha=2$ and $\beta=2$. The weights for the LLM Reasoning Inspired loss are set to $3$, $2$, and $1.5$ for the artifact, overexposed, and underexposed components, respectively. We conducted ablation experiments on these weights considering values from $1$ to $5$ and found that the selected combination of weights yield the best PSNR.